# Block Encoding of Sparse Matrices via Coherent Permutation

### v1

@@ 1 Introduction
Block encoding has emerged as a central primitive in modern quantum algorithms, providing a systematic way to embed a matrix into a unitary operator and thereby enabling polynomial transformation of operators via Quantum Singular Value Transformation (QSVT) [1]. The idea was first used implicitly in early breakthroughs such as the Harrow-Hassidim-Lloyd (HHL) algorithm for solving linear systems of equations [2] and Hamiltonian simulation techniques [3, 4, 5], and was later formalized by Gilyén et al. [1]. Since then, block encoding has become indispensible in a wide range of domains including quantum linear algebra, optimization, machine learning and quantum chemistry [6, 7, 8]. Succinctly, block encoding is the process of embedding a given (possibly non-unitary) matrix $A$ into a larger unitary operator $U_{A}$ as,
@@
Understanding the importance of block encoding has led to substantial research efforts to optimize its construction for different matrix classes. For arbitrary dense matrices, resource requirements have been well studied [9, 10]. Approximate block encodings using single- and two-qubit gates have been developed through the FABLE method [11, 12], and subsequent improvements using demultiplexor operations have been proposed [13]. For sparse matrices, block encodings have typically been formulated in terms of black-box oracles [1], but these works often omit explicit circuit-level implementations. More detailed realizations have been provided for structured sparsity [14]. Since the subnormalization factor $\alpha$ directly affects amplitude scaling in block encoding and consequently increases circuit depth in certain algorithms, recent works have sought to reduce it. Sünderhauf et al. [15] proposed schemes for matrices with arithmetic structure and PREP/UNPREP operators inspired by the Linear Combinations of Unitaries (LCU) method [5], while Yang et al. [16] introduced a dictionary-based protocol with improved subnormalization factor $\sum_{p=0}^{s-1}|A_{p}|$, where $s$ is the number of unique diagonal elements in the matrix. Despite these advances, efficient, and fully explicit constructions for block encodings of sparse matrices remain largely unexplored.
In this work, we introduce a novel quantum data manipulation strategy within the block encoding framework, providing enhanced control over matrix-space reachability while significantly reducing reliance on multi-controlled X (MCX) gates. Our method generalizes naturally and advances efficient quantum circuit mapping and compilation [17, 18, 19, 20], as well as reduces MCX control complexity [21, 22, 23]. We establish a novel connection between amplitude reordering by coherent permutation and combinatorial optimization, systematically assigning control qubits for MCX gates. This is particularly relevant for hardware with nearest-neighbor connectivity, where long-distance interactions increases noise and depth [24, 25, 26]. By optimizing control placement, our approach reduces circuit depth, improves hardware compatibility, and demonstrates practical benefits in block encoding of representative sparse matrices. We demonstrate our methods on structured sparse matrices, illustrating how theoretical advances translate into practical circuit implementations directly applicable to quantum algorithms.
@@ 2 Notations and Preprocessing
  • The Hamming distance [27, 28] between two binary strings $\vec{x}=(x_{0},x_{1},\cdots,x_{n-1})^{T}$ and $\vec{y}=(y_{0},y_{1},\cdots,y_{n-1})^{T}$, where $x_{k},y_{k}\in\{0,1\}$, is defined as
@@ 3 Block Encoding
In this section, we extend the PREP/UNPREP-based block encoding scheme to handle complex sparse matrices and present a clear, intuitive framework for constructing the corresponding quantum circuits. Building on the dictionary-based block encoding protocol [16], we define the $p^{\text{th}}$ data item $A_{p}$ from $A_{\text{data}}$, where $p\in[0,s-1]$, such that each matrix element $A_{ij}$ can be represented at row index $i$ and column index $j$. The set of all column indices associated with $A_{p}$ is denoted as $S_{c}(p)$. The corresponding row index is determined by $i=c_{p}(j)$, where $c_{p}(j)$ is an injective function mapping column indices $j$ to their respective row indices $i$.
@@
With these oracles in place, the block encoding scheme is formulated as described in Theorem 1 [16].
@@
Figure 3: 3(a) Circuit illustration of a data item $(a)$ from $m$ data qubits being selected and shifted by one column in an $8\times 8$ matrix formed by three matrix qubits $(|j\rangle)$. The control values for the left shift $(L_{a1})$ are $t_{1}=1$, while for the right shift $(R_{a1})$, they are $t_{1}=0$. The binary vector $\vec{a}$ (as defined in Eq. 2) is placed above the circle with a slash to indicate the control values according to the qubit reordering in Fig. 1. 3(b) Alternative circuit for shifting the same data item $(a)$, where the left shift $(L_{a1})$ has $t_{2}=0$ and the right shift $(R_{a1})$ has $t_{2}=1$. 3(c) Circuit representation of shifting data item $(b)$ by two columns $(L_{b2})$ and data item $(d)$ by four columns $(L_{d4})$ to the left in the same $8\times 8$ matrix. 3(d) A data item $(a)$ in the $r^{\text{th}}$ row $(L_{a}^{r}/R_{a}^{r})$ can be deleted by applying the control $(\vec{a})$ on the data qubits $(|\text{data}\rangle)$ and $(\vec{r})$ on the matrix qubits $(|j\rangle)$, followed by a NOT operation on the delete qubit $(|\text{del}\rangle)$.
@@ 3.1 State Preparation Oracle
The task of state preparation oracles PREP/UNPREP is to embed the data into amplitudes of a quantum state. Möttönen et al. [29, 30] introduced a state preparation method based on uniformly controlled rotation gates, which can be decomposed into either multi-controlled rotations or sequences of single- and two-qubit gates. This construction leverages classical preprocessing (e.g., Gray code ordering) to compute the required rotation angles efficiently. Later, Iten et al. [31] proposed a hardware-oriented approach that decomposes arbitrary isometries into single-qubit and CNOT gates, thereby reducing the CNOT count in practice. Both methods are exact, require no ancilla qubits, but generally scale exponentially in gate count and depth for arbitrary state preparation. More recent work has explored depth-optimized schemes [32], achieving circuit depths of $\mathcal{O}(m)$ for an $m$-qubit state or $\mathcal{O}(\log(ms))$ for $s$-sparse states. These improvements, however, often come at the cost of introducing additional ancilla qubits, which in some cases can scale exponentially with the problem size.
@@ 3.2 Index Mapping Oracle
The role of the oracle $O_{c}$ is to redistribute these values to their target positions within the matrix. This process involves two main steps: shifting [14] and deletion [15].
@@ 5 Combinatorial Optimization Based Mapping of Basis States
This is a linear assignment problem in its purest form and can be solved in $O(|S_{2}^{\prime}|^{3})$ time using the Hungarian method [33, 34, 35, 36]. The existence of an optimal solution to this problem establishes the existence of the required bijection $\phi$. Choosing the most common bit pattern $\tilde{a}$ Eq. 12 maximizes the overlap $|S_{2}\cap S_{3}|$ and therefore minimizes the residual set size $|S_{2}^{\prime}|$, which reduces the scale of the subsequent assignment problem and is advantageous for optimization.
In general, there may be multiple optimal bijections $\phi$ that achieve the same minimum total cost. One may then impose deterministic tie-breaking rules, such as lexicographic ordering of index pairs $(j,k)$. Alternatively, uniqueness can be enforced by perturbing the cost matrix, e.g. replacing each entry $C_{jk}$ with $C_{jk}+\epsilon r_{jk}$, where $\epsilon>0$ is arbitrarily small and $\{r_{jk}\}$ are distinct perturbations (random or deterministic). With such an $\epsilon$-perturbation, the minimizer is unique with probability one and corresponds to a canonical tie-broken solution of the original unperturbed problem [35, 37, 38].
@@ 7.2 Delete
Solution: The set of control states $S_{2}=\{0,1,4,7\}$ does not satisfy the constraint in Theorem 2, and therefore a permutation is required, as illustrated in Fig. 5(d). Following Theorem 3, the bijection $\phi:S_{2}\mapsto S_{3}$ is obtained using the linear sum assignment algorithm [39], yielding $\phi:S_{2}=\{0,1,4,7\}\mapsto S_{3}=\{0,1,2,3\}$. The corresponding permutation of amplitudes is then implemented using the gate constructions described in Section 6:
@@ 9 Applications
In this section, we present two examples of sparse matrix block encoding: a tridiagonal complex matrix and a structured real matrix.
9.1 Tridiagonal Complex Matrix
Consider a tridiagonal complex matrix $A\in\mathbb{C}^{2^{n}\times 2^{n}}$ of the form,
@@ 9.1 Tridiagonal Complex Matrix
Figure 7: Circuit representation for block encoding tridiagonal complex matrix Eq. 23. The control states are represented by binary representation as in Fig. 3.
@@ 9.2 Structured Real Matrix
The next step is to determine whether amplitude permutation is required. For demonstration purposes, we consider the shift $L_{1}$ in Table 2 and the deletion $D_{a1}$ in Table 1. For $L_{1}$, the control states $S_{2}$ corresponding to the data items $a_{0},a_{1},a_{7},a_{8},a_{11},a_{12},a_{14},a_{15}]$ do not satisfy the constraint in Theorem 2. Therefore, we apply the permutation operator $\text{A\_PERMUTE}_{\text{R}}$ as shown in Fig. 5(a). Following Theorem 3, the mapping $\phi:S_{2}\mapsto S_{3}$ is obtained using the linear sum assignment algorithm [39]. Finally, the gates are generated to permute $(S_{2},S_{3})$ according to Definition 6.2, resulting in the following transformation:
@@ 10 Discussion
Finally we validated our framework by implementing it on two respective cases: a tridiagonal complex matrix and a structured real matrix, demonstrating that the entire pipeline from theory to circuit operates seamlessly. The resulting circuits are immediately applicable to high-impact quantum algorithms such as QSVT, HHL, and Hamiltonian simulation [40], thereby extending the reach of block encoding into practical algorithm design and realistic hardware execution.

### v2

@@ 1 Introduction
Block encoding has emerged as a central primitive in modern quantum algorithms, providing a systematic way to embed a matrix into a unitary operator and thereby enabling polynomial transformation of operators via Quantum Singular Value Transformation (QSVT) gilyen2019quantum . The idea was first used implicitly in early breakthroughs such as the Harrow-Hassidim-Lloyd (HHL) algorithm for solving linear systems of equations harrow2009quantum and Hamiltonian simulation techniques berry2015simulating , childs2017quantum , childs2012hamiltonian , and was later formalized by Gilyén et al. gilyen2019quantum . Since then, block encoding has become indispensible in a wide range of domains including quantum linear algebra, optimization, machine learning and quantum chemistry brandao2017quantum , van2017quantum , babbush2018encoding . Succinctly, block encoding is the process of embedding a given (possibly non-unitary) matrix $A$ into a larger unitary operator $U_{A}$ as,
@@
Understanding the importance of block encoding has led to substantial research efforts to optimize its construction for different matrix classes. For arbitrary dense matrices, resource requirements have been well studied clader2023quantum , chakraborty2018power . Approximate block encodings using single- and two-qubit gates have been developed through the FABLE method camps2022fable , kuklinski2024s , and subsequent improvements using demultiplexor operations have been proposed li2025binary . For sparse matrices, block encodings have typically been formulated in terms of black-box oracles gilyen2019quantum , but these works often omit explicit circuit-level implementations. More detailed realizations have been provided for structured sparsity camps2024explicit . Since the subnormalization factor $\alpha$ directly affects amplitude scaling in block encoding and consequently increases circuit depth in certain algorithms, recent works have sought to reduce it. Sünderhauf et al. sunderhauf2024block proposed schemes for matrices with arithmetic structure and PREP/UNPREP operators inspired by the Linear Combinations of Unitaries (LCU) method childs2012hamiltonian , while Yang et al. yang2024block introduced a dictionary-based protocol with improved subnormalization factor $\sum_{p=0}^{s-1}|A_{p}|$, where $s$ is the number of unique diagonal elements in the matrix. Despite these advances, efficient, and fully explicit constructions for block encodings of sparse matrices remain largely unexplored.
In this work, we introduce a novel quantum data manipulation strategy within the block encoding framework, providing enhanced control over matrix-space reachability while significantly reducing reliance on multi-controlled X (MCX) gates. Our method generalizes naturally and advances efficient quantum circuit mapping and compilation shende2005synthesis , amy2013meet , cowtan2019qubit , murali2019noise , as well as reduces MCX control complexity barenco1995elementary , miller2003transformation , maslov2016advantages . We establish a novel connection between amplitude reordering by coherent permutation and combinatorial optimization, systematically assigning control qubits for MCX gates. This is particularly relevant for hardware with nearest-neighbor connectivity, where long-distance interactions increases noise and depth linke2017experimental , beals2013efficient , kutin2007computation . By optimizing control placement, our approach reduces circuit depth, improves hardware compatibility, and demonstrates practical benefits in block encoding of representative sparse matrices. We demonstrate our methods on structured sparse matrices, illustrating how theoretical advances translate into practical circuit implementations directly applicable to quantum algorithms.
@@ 2 Notations and Preprocessing
  • The Hamming distance hamming1950error , li2022quantum between two binary strings $\vec{x}=(x_{0},x_{1},\cdots,x_{n-1})^{T}$ and $\vec{y}=(y_{0},y_{1},\cdots,y_{n-1})^{T}$, where $x_{k},y_{k}\in\{0,1\}$, is defined as
@@ 3 Block Encoding
In this section, we extend the PREP/UNPREP-based block encoding scheme to handle complex sparse matrices and present a clear, intuitive framework for constructing the corresponding quantum circuits. Building on the dictionary-based block encoding protocol yang2024block , we define the $p^{\text{th}}$ data item $A_{p}$ from $A_{\text{data}}$, where $p\in[0,s-1]$, such that each matrix element $A_{ij}$ can be represented at row index $i$ and column index $j$. The set of all column indices associated with $A_{p}$ is denoted as $S_{c}(p)$. The corresponding row index is determined by $i=c_{p}(j)$, where $c_{p}(j)$ is an injective function mapping column indices $j$ to their respective row indices $i$.
@@
With these oracles in place, the block encoding scheme is formulated as described in Theorem 1 yang2024block .
@@
Figure 3: 3(a) Circuit illustration of a data item $(a)$ from $m$ data qubits being selected and shifted by one column in an $8\times 8$ matrix formed by three matrix qubits $(|j\rangle)$. The control values for the left shift $(L_{a1})$ are $t_{1}=1$, while for the right shift $(R_{a1})$, they are $t_{1}=0$. The binary vector $\vec{a}$ (as defined in Eq. 2) is placed above the circle with a slash to indicate the control values according to the qubit ordering in Fig. 1. 3(b) Alternative circuit for shifting the same data item $(a)$, where the left shift $(L_{a1})$ has $t_{2}=0$ and the right shift $(R_{a1})$ has $t_{2}=1$. 3(c) Circuit representation of shifting data item $(b)$ by two columns $(L_{b2})$ and data item $(d)$ by four columns $(L_{d4})$ to the left in the same $8\times 8$ matrix. 3(d) A data item $(a)$ in the $r^{\text{th}}$ row $(L_{a}^{r}/R_{a}^{r})$ can be deleted by applying the control $(\vec{a})$ on the data qubits $(|\text{data}\rangle)$ and $(\vec{r})$ on the matrix qubits $(|j\rangle)$, followed by a NOT operation on the delete qubit $(|\text{del}\rangle)$.
@@ 3.1 State Preparation Oracle
The task of state preparation oracles PREP/UNPREP is to embed the data into amplitudes of a quantum state. Möttönen et al. mottonen2004transformation , mottonen2004quantum introduced a state preparation method based on uniformly controlled rotation gates, which can be decomposed into either multi-controlled rotations or sequences of single- and two-qubit gates. This construction leverages classical preprocessing (e.g., Gray code ordering) to compute the required rotation angles efficiently. Later, Iten et al. iten2016quantum proposed a hardware-oriented approach that decomposes arbitrary isometries into single-qubit and CNOT gates, thereby reducing the CNOT count in practice. Both methods are exact, require no ancilla qubits, but generally scale exponentially in gate count and depth for arbitrary state preparation. More recent work has explored depth-optimized schemes zhang2022quantum , achieving circuit depths of $\mathcal{O}(m)$ for an $m$-qubit state or $\mathcal{O}(\log(ms))$ for $s$-sparse states. These improvements, however, often come at the cost of introducing additional ancilla qubits, which in some cases can scale exponentially with the problem size.
@@ 3.2 Index Mapping Oracle
The role of the oracle $O_{c}$ is to redistribute these values to their target positions within the matrix. This process involves two main steps: shifting camps2024explicit and deletion sunderhauf2024block .
@@ 5 Combinatorial Optimization Based Mapping of Basis States
This is a linear assignment problem in its purest form and can be solved in $O(|S_{2}^{\prime}|^{3})$ time using the Hungarian method kuhn1955hungarian , munkres1957algorithms , burkard2012assignment , wolsey2020integer . The existence of an optimal solution to this problem establishes the existence of the required bijection $\phi$. Choosing the most common bit pattern $\tilde{a}$ Eq. 12 maximizes the overlap $|S_{2}\cap S_{3}|$ and therefore minimizes the residual set size $|S_{2}^{\prime}|$, which reduces the scale of the subsequent assignment problem and is advantageous for optimization.
In general, there may be multiple optimal bijections $\phi$ that achieve the same minimum total cost. One may then impose deterministic tie-breaking rules, such as lexicographic ordering of index pairs $(j,k)$. Alternatively, uniqueness can be enforced by perturbing the cost matrix, e.g. replacing each entry $C_{jk}$ with $C_{jk}+\epsilon r_{jk}$, where $\epsilon>0$ is arbitrarily small and $\{r_{jk}\}$ are distinct perturbations (random or deterministic). With such an $\epsilon$-perturbation, the minimizer is unique with probability one and corresponds to a canonical tie-broken solution of the original unperturbed problem burkard2012assignment , schrijver2003combinatorial , korte2008combinatorial .
@@ 7.2 Delete
Solution: The set of control states $S_{2}=\{0,1,4,7\}$ does not satisfy the constraint in Theorem 2, and therefore a permutation is required, as illustrated in Fig. 5(d). Following Theorem 3, the bijection $\phi:S_{2}\mapsto S_{3}$ is obtained using the linear sum assignment algorithm crouse2016implementing , yielding $\phi:S_{2}=\{0,1,4,7\}\mapsto S_{3}=\{0,1,2,3\}$. The corresponding permutation of amplitudes is then implemented using the gate constructions described in Section 6:
@@ 9 Applications
In this section, we present two examples of sparse matrix block encoding: a complex tridiagonal matrix and a structured real matrix.
9.1 Complex Tridiagonal Matrix
Consider a complex tridiagonal matrix $A\in\mathbb{C}^{2^{n}\times 2^{n}}$ of the form,
@@ 9.1 Complex Tridiagonal Matrix
Figure 7: Circuit representation for block encoding complex tridiagonal matrix Eq. 23. The control states are represented by binary representation as in Fig. 3.
@@ 9.2 Structured Real Matrix
The next step is to determine whether amplitude permutation is required. For demonstration purposes, we consider the shift $L_{1}$ in Table 2 and the deletion $D_{a1}$ in Table 1. For $L_{1}$, the control states $S_{2}$ corresponding to the data items $a_{0},a_{1},a_{7},a_{8},a_{11},a_{12},a_{14},a_{15}]$ do not satisfy the constraint in Theorem 2. Therefore, we apply the permutation operator $\text{A\_PERMUTE}_{\text{R}}$ as shown in Fig. 5(a). Following Theorem 3, the mapping $\phi:S_{2}\mapsto S_{3}$ is obtained using the linear sum assignment algorithm crouse2016implementing . Finally, the gates are generated to permute $(S_{2},S_{3})$ according to Definition 6.2, resulting in the following transformation:
@@ 10 Discussion
Finally we validated our framework by implementing it on two respective cases: a complex tridiagonal matrix and a structured real matrix, demonstrating that the entire pipeline from theory to circuit operates seamlessly. The resulting circuits are immediately applicable to high-impact quantum algorithms such as QSVT, HHL, and Hamiltonian simulation martyn2021grand , thereby extending the reach of block encoding into practical algorithm design and realistic hardware execution.